# Distributionally Robust Receive Beamforming

### v2

@@ VII Conclusions
This article introduces a unified mathematical framework for receive beamforming of wireless signals from the perspective of data-driven machine learning, which reveals that channel estimation is not a necessary operation. To combat the limited pilot size and several types of uncertainties in the signal model, the distributionally robust (DR) receive beamforming framework is then suggested. We prove that the diagonal-loading (DL) methods are distributionally robust against the scarcity of pilot data and the uncertainties in the signal model. In addition, we generalize the diagonal-loading methods to achieve better estimation performance (e.g., the DR Wiener beamformer using $F$-norm for uncertainty quantification), at the cost of significantly higher computational burdens. Experiments suggest that nonlinear beamformers such as the Kernel and the Kernel-DL methods have the potential when the pilot size is small and/or the signal model is not linear Gaussian. Compared with the Kernel and the Kernel-DL beamformers, neural-network-based solutions [10, 12] have the stronger expressive capability of nonlinearities, which however are unscalable in the numbers of transmit and receive antennas, and significantly more time-consuming in training and more troublesome in tuning hyper-parameters (e.g., the number of layers and the number of neurons in each layer) than the studied eleven beamformers.
Supplementary Materials
@@ Appendix B Details on Real-Space Signal Representation
Let $\bm{C}_{x}\coloneqq\mathbb{E}\mathbf{x}\mathbf{x}^{\mathsf{T}}$, $\bm{C}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{T}}$, and $\bm{C}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{T}}=\bm{0}$. We have
@@ C-B Non-Stationary Channel Statistics
In the main body of the paper (see also Fact 1), we assume that the true data-generating distribution $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ is time-invariant within a frame. In real-world operations, however, this assumption might be untenable.
@@
Let $\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ denote the true data-generating distribution at time point $t_{i}$ where $i=0,1,2,\ldots,8$. Specifically, we have $(\mathbf{x}_{i},\mathbf{s}_{i})\sim\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ for every $i$. Therefore, the pilot data set (i.e., the training data set) $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),(\bm{x}_{3},\bm{s}_{3}),(\bm{x}_{4},\bm{s}_{4})\}$ can be seen as realizations of the mean distribution $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}$ of underlying true training-data distributions where $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\sum^{4}_{i=1}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, which is a mixture distribution with mixing weights $0\leq h_{1},h_{2},h_{3},h_{4}\leq 1$; $\sum^{4}_{i=1}h_{i}=1$. Similarly, the communication data set (i.e., the testing data set) $\{(\bm{x}_{5},\bm{s}_{5}),(\bm{x}_{6},\bm{s}_{6}),(\bm{x}_{7},\bm{s}_{7}),(\bm{x}_{8},\bm{s}_{8})\}$ can be seen as realizations of the mean $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$ of the underlying true testing-data distributions where $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}=\sum^{8}_{i=5}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, with mixing weights $0\leq h_{5},h_{6},h_{7},h_{8}\leq 1$; $\sum^{8}_{i=5}h_{i}=1$.
@@
Therefore, the distributionally robust modeling and solution framework is still valid to hedge against the distributional uncertainty in the nominal distribution ^$\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}}$ compared to the underlying true distribution $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$. When $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, as assumed in the main body of the paper, we have $\epsilon_{1}\to 0$ and $\epsilon\to\epsilon_{2}=0$ as the pilot size tends to infinity; however, when $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}\neq\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, the radius $\epsilon\to\epsilon_{2}\neq 0$ although $\epsilon_{1}\to 0$.
@@
$d(\mathbb{P}_{0,\mathbf{x},\mathbf{s},i},~{}\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}})\leq\epsilon,~{}~{}~{}\forall i\in\{0,1,2,\ldots,8\}.$
@@
Appendix D Proof of Lemma 1
@@
where $\langle\bm{A},\bm{B}\rangle\coloneqq\operatorname{Tr}\bm{A}^{\mathsf{H}}\bm{B}$ for two matrices $\bm{A}$ and $\bm{B}$. Therefore, the objective function of (19) is convex in $\bm{W}$ and linear (thus concave) in the matrix variable $\bm{R}$. Note also that complex Euclidean spaces are linear topological spaces. Hence, due to Sion’s minimax theorem, Problem (19) is equivalent to
@@
Appendix E Proof of Theorem 1
@@
$[\bm{x}^{\mathsf{H}},~{}\bm{y}^{\mathsf{H}}]\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right]\left[\begin{array}[]{c}\bm{x}\\ \bm{y}\end{array}\right]=\|\bm{W}\bm{x}-\bm{y}\|^{2}_{F}\geq 0.$
@@
Appendix F Proof of Proposition 3
@@
$\min_{\bm{W}\in\mathbb{R}^{2M\times 2N}}\max_{\mathbb{P}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}\in\mathcal{U}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}[\bm{W}\underline{\mathbf{z}}-\underline{\mathbf{s}}][\bm{W}\underline{\mathbf{z}}-\underline{\mathbf{s}}]^{\mathsf{T}}.$ (80)
@@
Appendix G Additional Experimental Results

### v3

@@ VII Conclusions
This article introduces a unified mathematical framework for receive combining of wireless signals from the perspective of data-driven machine learning, which reveals that channel estimation is not a necessary operation. To combat the limited pilot size and several types of uncertainties in the signal model, the distributionally robust (DR) receive combining framework is then suggested. We prove that the diagonal-loading (DL) methods are distributionally robust against the scarcity of pilot data and the uncertainties in the signal model. In addition, we generalize the diagonal-loading methods to achieve better estimation performance (e.g., the DR Wiener beamformer using $F$-norm for uncertainty quantification), at the cost of significantly higher computational burdens. Experiments suggest that nonlinear combiners such as the Kernel and the Kernel-DL methods have the potential when the pilot size is small and/or the signal model is not linear Gaussian. Compared with the Kernel and the Kernel-DL combiners, neural-network-based solutions [10, 12] have a stronger expressive capability of nonlinearities, which however are unscalable in the numbers of transmit and receive antennas, and significantly more time-consuming in training and more troublesome in tuning hyper-parameters (e.g., the number of layers and the number of neurons in each layer) than the studied eleven combiners.
@@ Appendix B Details on Real-Space Signal Representation
Let $\bm{R}_{x}\coloneqq\mathbb{E}\mathbf{x}\mathbf{x}^{\mathsf{H}}$, $\bm{C}_{x}\coloneqq\mathbb{E}\mathbf{x}\mathbf{x}^{\mathsf{T}}$, $\bm{C}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{T}}$, and $\bm{C}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{T}}=\bm{0}$. We have
@@ C-B Non-Stationary Channel Statistics
In the main body of the article (see also Fact 1), we assume that the true data-generating distribution $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ is time-invariant within a frame. In real-world operations, however, this assumption might be untenable.
@@
Let $\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ denote the true data-generating distribution at time point $t_{i}$ where $i=1,2,\ldots,8$. Specifically, we have $(\mathbf{x}_{i},\mathbf{s}_{i})\sim\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$ for every $i$. Therefore, the pilot data set (i.e., the training data set) $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),(\bm{x}_{3},\bm{s}_{3}),(\bm{x}_{4},\bm{s}_{4})\}$ can be seen as realizations of the mean distribution $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}$ of underlying true training-data distributions where $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\sum^{4}_{i=1}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, which is a mixture distribution with mixing weights $0\leq h_{1},h_{2},h_{3},h_{4}\leq 1$; $\sum^{4}_{i=1}h_{i}=1$. Similarly, the communication data set (i.e., the testing data set) $\{(\bm{x}_{5},\bm{s}_{5}),(\bm{x}_{6},\bm{s}_{6}),(\bm{x}_{7},\bm{s}_{7}),(\bm{x}_{8},\bm{s}_{8})\}$ can be seen as realizations of the mean $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$ of the underlying true testing-data distributions where $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}=\sum^{8}_{i=5}h_{i}\mathbb{P}_{0,\mathbf{x},\mathbf{s},i}$, with mixing weights $0\leq h_{5},h_{6},h_{7},h_{8}\leq 1$; $\sum^{8}_{i=5}h_{i}=1$.
@@
Therefore, the distributionally robust modeling and solution framework is still valid to hedge against the distributional uncertainty in the nominal distribution ^$\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}}$ compared to the underlying true distribution $\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$. When $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}=\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, as assumed in the main body of the article, we have $\epsilon_{1}\to 0$ and $\epsilon\to\epsilon_{2}=0$ as the pilot size tends to infinity; however, when $\mathbb{P}_{\text{train},0,\mathbf{x},\mathbf{s}}\neq\mathbb{P}_{\text{test},0,\mathbf{x},\mathbf{s}}$, the radius $\epsilon\to\epsilon_{2}\neq 0$ although $\epsilon_{1}\to 0$.
@@
$d(\mathbb{P}_{0,\mathbf{x},\mathbf{s},i},~{}\hat{\mathbb{P}}_{\text{train},\mathbf{x},\mathbf{s}})\leq\epsilon,~{}~{}~{}\forall i\in\{1,2,\ldots,8\}.$
@@
Appendix D Additional Discussions on Distributionally Robust Estimation
To develop this article, the typical minimum mean-squared error (MSE) criterion is employed; see (7) and (10). Accordingly, the distributionally robust receive combining framework in this article is exemplified using the MSE cost function. The cost function for wireless signal estimation, however, can be any Borel-measurable function $h:\mathbb{C}^{M}\times\mathbb{C}^{M}\to\mathbb{R}_{+}$. As a result, the optimal estimation problem under the distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is given by
$\min_{\bm{\phi}\in\mathcal{B}_{\mathbb{C}^{N}\to\mathbb{C}^{M}}}\mathbb{E}_{\mathbf{x},\mathbf{s}}h[\bm{\phi}(\mathbf{x}),\mathbf{s}].$ (76)
Specific examples of $h$ in wireless communications can be, e.g., mean absolute error, Huber’s cost function [37, 38] where $h$ is no longer quadratic as in (7) and (10). Accordingly, when the distributional uncertainty exists in $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, the distributionally robust receive combining framework becomes
$\min_{\bm{\phi}\in\mathcal{B}_{\mathbb{C}^{N}\to\mathbb{C}^{M}}}\max_{\mathbb{P}_{\mathbf{x},\mathbf{s}}\in\mathcal{U}_{\mathbf{x},\mathbf{s}}}\mathbb{E}_{\mathbf{x},\mathbf{s}}h[\bm{\phi}(\mathbf{x}),\mathbf{s}].$ (77)
Problem (77) is generally challenging to solve because it is an infinite-dimensional program. Therefore, in practice, we can limit the feasible region of $\bm{\phi}$ to a parameterized subspace of $\mathcal{B}_{\mathbb{C}^{N}\to\mathbb{C}^{M}}$, for example, a reproducing kernel Hilbert space $\mathcal{H}$ or a neural network function space $\mathcal{K}$; see Section II. Consequently, Problem (77) is approximated by the following finite-dimensional (in terms of $\bm{W}$) program
$\min_{\bm{W}}\max_{\mathbb{P}_{\mathbf{x},\mathbf{s}}\in\mathcal{U}_{\mathbf{x},\mathbf{s}}}\mathbb{E}_{\mathbf{x},\mathbf{s}}h[\bm{\phi}_{\bm{W}}(\mathbf{x}),\mathbf{s}],$ (78)
where $\bm{W}$ parameterizes $\bm{\phi}$ and lies in real or complex coordinate spaces; note that both $\mathcal{H}$ and $\mathcal{K}$ can be dense in $\mathcal{B}$. Under the MSE cost function, (78) is particularized in (19) for linear function spaces, in (57) for reproducing kernel Hilbert spaces, and in (69) for neural network function spaces, which build this article in a technically tractable manner.
The distributionally robust receive combining problem (77) under generic cost functions $h$ and generic feasible regions of $\bm{\phi}$ can be technically challenging. Even for the simplified problem (78), the solution method can be quite complex, and closed-form solutions cannot be generally guaranteed; see, e.g., [39, 40]. The complication further arises when the distributional uncertainty sets $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ for $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ are complicated; see, e.g., [32]. Therefore, this article serves as the starting point of distributionally robust receive combining, in which closed-form solutions are largely ensured by leveraging
  1.    F1)
    the MSE cost function as in (7) and (10);
  2.    F2)
    the linear function spaces as in (19) and reproducing kernel Hilbert spaces as in (57);
  3.    F3)
    the second-moment-based uncertainty sets in Definitions 1, 2, 3, and 4; see also Corollary 1, Claim 2, and Example 4.
Note that even under the features F1) and F2), the closed-form solutions cannot be guaranteed. For example, if Wasserstein or F-norm uncertainty sets are used, the associated distributionally robust receive combining problems can be computationally heavy; see (47) and (52) as well as Propositions 1 and 2. However, for emerging high-performance computing devices, the computational burden may be no longer an issue in the future. Hence, advanced distributionally robust receive combining formulations based on (77) and (78) are still attractive for future-generation communication systems. This article seeks to provide a foundation for this direction.
Appendix E Proof of Lemma 1
@@
where $\langle\bm{A},\bm{B}\rangle\coloneqq\operatorname{Tr}\bm{A}^{\mathsf{H}}\bm{B}$ for two matrices $\bm{A}$ and $\bm{B}$. Therefore, the objective function of (19) is convex in $\bm{W}$ and linear (thus concave) in the matrix variable $\bm{R}$. Hence, due to Sion’s minimax theorem [41, Corollary 3.3], Problem (19) is equivalent to
@@
Note that the feasible region of $\bm{R}$ is compact convex, and that of $\bm{W}$ (i.e., $\mathbb{C}^{M\times N}$) is convex.
@@
Appendix F Proof of Theorem 1
@@
$[\bm{x}^{\mathsf{H}},~{}\bm{y}^{\mathsf{H}}]\left[\begin{array}[]{cc}\bm{W}^{\mathsf{H}}\bm{W}&-\bm{W}^{\mathsf{H}}\\ -\bm{W}&\bm{I}_{M}\end{array}\right]\left[\begin{array}[]{c}\bm{x}\\ \bm{y}\end{array}\right]=\|\bm{W}\bm{x}-\bm{y}\|^{2}_{2}\geq 0.$
@@
Appendix G Proof of Proposition 3
@@
$\min_{\bm{W}\in\mathbb{R}^{2M\times L}}\max_{\mathbb{P}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}\in\mathcal{U}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{z}},\underline{\mathbf{s}}}[\bm{W}\underline{\mathbf{z}}-\underline{\mathbf{s}}][\bm{W}\underline{\mathbf{z}}-\underline{\mathbf{s}}]^{\mathsf{T}}.$ (83)
@@
Supplementary Materials
Appendix H Additional Experimental Results